### v1

# Regular Kerr black holes: The fellowship of the string

Abstract
In their interesting article entitled "Kerrr" black hole: The lord of the string [Phys. Lett. B 688, 82 (2010)], Smailagic and Spallucci report on a regular Kerr-like black hole spacetime with a string in the central region. The metric is of the Kerr-Schild form with a particular mass function, that is chosen in order to avoid the ring singularity of the Kerr metric. However, as it is known, even for a class of well-behaved mass functions, the curvature scalars present a discontinuity in the equatorial plane at the ring. The authors associate such a discontinuity with the presence of a string of matter and join the interior and exterior regions of the string by making a particular choice of the orientations of the normal vectors at each side of the ring. Differently from what was interpreted by the authors, we show here that the string is not rotating in the scenario of that work. We also show that the choice regarding the normal vector orientation leads to a geometry that is not compatible with the center of a rotating black hole in an asymptotically flat spacetime. We then consider all the four possible combinations of the normal vector orientation at each side of the ring and, by using the Darmois-Israel formalism, interpret each resulting solution. We show that, out of the four possibilities, only one of the four models for the thin shell solution at the ring yields the appropriate geometry. More importantly, this solution does not correspond to the one obtained by Smailagic and Spallucci. Finally, we also discuss a scenario where the string is indeed rotating.
Regular black holes; Kerr-like geometries ; thin-shell formalism
I Introduction
The investigation of regular rotating black holes is motivated by the complications within the interior of the Kerr geometry [1], which, while providing a realistic description of the exterior of a rotating black hole, presents issues such as a ring singularity, causality violations, and closed timelike curves within the event horizon [2]. A potential remedy involves replacing the problematic interior region with a regular matter source, mirroring the approach taken for static regular black holes.
In Ref. [3], the authors demonstrated that, in the context of Kerr-Schild spacetimes [4], a Lorentzian coordinate system can be selected in such a way that it leads to the linearization of the Einstein field equations. As an application of this result, the authors showed that complex translations can be applied in order to obtain new solutions for spinning systems, including new anisotropic interior metrics that may be matched to the Kerr metric on an oblate spheroid. This procedure paved the way for constructing rotating counterparts of static and spherically symmetric regular black hole solutions resembling the Kerr metric in the asymptotic region, but with the constant mass parameter replaced by a mass function $m(r)$, so modifying the Kerr geometry close to the central ring. Subsequent explorations have delved into understanding the source generating the rotating metric [5, 6] and a substantial body of work has emerged on regular rotating objects [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 22, 21, 23, 24, 25].
In contrast to the Kerr metric, it was shown by Torres [15] and Maeda [21] that, if the mass function $m(r)$ can be expanded around $r=0$ as $m(r)\approx m_{0}r^{3+\alpha}$ with $m_{0}\neq 0$ and $\alpha\geq 0$, the ring $\mathcal{S}\!\!:\!(r=0,\,\theta=\pi/2)$ corresponds to a conical singularity when extended to the region $r<0$, and not to a scalar polynomial curvature singularity. Moreover, if the extension to the region $r<0$ is not performed, Torres [26] showed that the ring is devoid of conical singularities as well.
Even though the different types of singularities can be avoided in the ring for the Gürses-Gürsey rotating metric, it is well known that the curvature scalars are not well defined there, once they present a finite jump which depends on the path taken when approaching the ring for $\alpha=0$ [7, 8, 15, 21, 19, 24, 25]. On the other hand, if $\alpha>0$ the curvature scalars are continuous [26]. This feature exhibited by the Gürses-Gürsey metric, for $\alpha=0$, has been interpreted by Smailagic and Spallucci [7] as a string substituting the ring singularity observed in the Kerr solution. More specifically, since the limit $r\to 0$ for $\theta=\pi/2$ corresponds to the curvature of a de Sitter-like geometry, while the limit $r\to 0$ for $\theta\neq\pi/2$ corresponds to the curvature of a Minkowski-like geometry [15], the authors glued a Minkowski-like spacetime with a de Sitter-like spacetime at the ring and interpreted the discontinuity in the extrinsic curvature as due to the presence of a string of matter for a unconventional choice of the signs of the normal vectors at ring.
In this work we explore all possible choices of the orientations of the normal vectors at ring. Let us remember that the normal vector of a thin shell plays an important role, since it the determines extrinsic curvature of the shell regarding the spacetime in which the shell is immersed [27]. In fact, to uncover all possible thin shell solutions, it is necessary to take into account that the normal to a shell can have two distinct orientations relative to the center of the coordinates of both the interior and exterior spacetimes (or spacetime regions), even though the normal vector is fixed to point always from a given spacetime to the other [28, 29]. Moreover, we also show that, in the scenario presented in Ref. [7], the string replacing the Kerr ring is actually static. To address this issue, we construct a scenario in which the string is indeed rotating.
II The regular rotating black hole
II.1 The spacetime metric
In Boyer-Lindquist coordinates, the metric of interest here reads
$ds^{2}=$ $-\left(1-\frac{2r\,m(r)}{\Sigma(r,\theta)}\right)dt^{2}+\frac{\Sigma(r,\theta)}{\Delta(r)}dr^{2}$ (1)
$+\Sigma(r,\theta)d\theta^{2}-\frac{4r\,m(r)a\sin^{2}\theta}{\Sigma}dt\,d\varphi$
$+\left(r^{2}+a^{2}+\frac{2r\,m(r)a^{2}\sin^{2}\theta}{\Sigma(r,\theta)}\right)\sin^{2}\theta\,d\varphi^{2},$
where $m(r)$ stands for the mass function, $a$ indicates the rotation parameter, and $\Sigma(r,\theta)$ and $\Delta(r)$ are defined by
$\Sigma(r,\theta)=r^{2}+a^{2}\cos^{2}\theta,$ (2)
$\Delta(r)=r^{2}+a^{2}-2rm(r),$ (3)
respectively.
In order to make our analysis as general as possible, we do not choose any specific mass function $m(r)$. The only imposition on $m(r)$ is that it can be expanded around $r=0$ as
$m(r)=m_{0}r^{3+\alpha},$ (4)
with $m_{0}\neq 0$ and $\alpha\geq 0$. This assumption guarantees that the desired properties of metric (1) as summarized in Sec. I remain valid, it assures particularly that the curvature scalars do not diverge in the limit $r\to 0$. This is the main reason why such an asymptotic behavior of the mass function is present in most of the works cited here, specially in the work by Smailagic and Spallucci [7].
The Ricci scalar for the metric (1) is given by
$\mathcal{R}=\frac{2}{\Sigma(r,\theta)}\big{[}2m^{\prime}(r)+r\,m^{\prime\prime}(r)\big{]},$ (5)
with the primes indicating differentiation with respect to the coordinate $r$. Equations (4) and (5) tell us that, around $r=0$, the Ricci scalar is given approximately by
$\mathcal{R}=2\left(3+\alpha\right)\left(4+\alpha\right)\frac{m_{0}\,r^{2+\alpha}}{\Sigma(r,\theta)}.$ (6)
Hence, for $\theta\neq\pi/2$, the limit $r\to 0$ gives $\mathcal{R}=0$ for all $\alpha\geq 0$. In turn, for $\theta=\pi/2$, Eq. (6) furnishes $2\left(3+\alpha\right)\left(4+\alpha\right)m_{0}r^{\alpha}$ and, therefore, the limit $r\to 0$ gives $\mathcal{R}=0$ for $\alpha>0$, but it gives $\mathcal{R}=24m_{0}$ for $\alpha=0$. In conclusion, if the mass function is of the form $m(r)\sim r^{3}$ close to $r\to 0$, then the Ricci scalar is not well defined there. The same also happens with the other independent curvatures scalars [15].
As just mentioned above, the singular locus of the Kerr metric parameterized by ($r=0,\,\theta=\pi/2$) in the Boyer-Lindquist coordinates can be regularized by replacing the constant mass parameter of the Kerr metric with a power-law mass function of the form (4), but just for $\alpha>0$. For $\alpha<0$ , the locus is still a curvature singularity as in the Kerr metric. The $\alpha=0$ case is special, since the curvature scalars are all finite but present a discontinuity at that locus. This feature motivated the authors of Ref. [7] to introduce a string replacing the ring singularity of the Kerr spacetime. Following this idea, in the present work we explore the proposal by Smailagic and Spallucci in more detail.
II.2 Inside the ring: The disk
Let us notice that the limit $r\to 0$ for $\theta\neq\pi/2$ in Eq. (1) corresponds to a disk $\mathcal{D}$ of radius $a$ that may be parameterized by the polar coordinates $(\rho,\,\varphi)$ such that $0\leq\rho\leq a$ and $0\leq\varphi\leq 2\pi$. The coordinate $\rho$ is given in terms of the Boyer-Lindquist coordinates by
$\rho=a\sin\theta,$ (7)
while $\varphi$ is the azimuth angle. At the disk $\mathcal{D}\!:(r=0,0\leq\theta<\pi/2)$, the metric (1) reduces to
$ds^{2}_{i}=-dt^{2}+d\rho^{2}+\rho^{2}d\varphi^{2},$ (8)
which is the flat metric, indicating a Minkowskian region with vanishing curvature scalars.
II.3 The ring
The ring $\mathcal{S}$ is at the boundary of the disk defined in the last paragraph, i.e., it may be parameterized by $(\rho=a,0\leq\varphi\leq 2\pi)$. At the ring $\mathcal{S}\!:(r=0,\theta=\pi/2)$, metric (1) reduces to
$ds^{2}_{\mathcal{S}}=-dt^{2}+a^{2}d\varphi^{2},$ (9)
where, from Eq. (7), $r=0$ and $\theta=\pi/2$ implies $\rho=a$.
II.4 Outside the ring: The equatorial plane
For the present analysis it is useful to have the explicit form of metric (1) at the equatorial plane, i.e., for $\theta=\pi/2$. First we define a new radial coordinate $\varrho$ by
$\varrho=\sqrt{r^{2}+a^{2}},$ (10)
whose range is $\varrho\in[a,\infty)$. With this, at $\theta=\pi/2$, metric (1) assumes the form
$ds^{2}_{\scriptscriptstyle{E}}=$ $-\left(1-\frac{2m(\varrho)}{\sqrt{\varrho^{2}-a^{2}}}\right)dt^{2}-\frac{4m(\varrho)a}{\sqrt{\varrho^{2}-a^{2}}}dt\,d\varphi$ (11)
$+\frac{\varrho^{2}}{\Delta(\varrho)}d\varrho^{2}+\left(\varrho^{2}+\frac{2m(\varrho)a^{2}}{\sqrt{\varrho^{2}-a^{2}}}\right)d\varphi^{2},$
where $\Delta(\varrho)=\varrho^{2}-2m(\varrho)\sqrt{\varrho^{2}-a^{2}}$.
Approaching the ring, but remaining outside, i.e, in the region $\mathcal{A}\!:\!(r\to 0,\,\theta=\pi/2)$, we have $m(\varrho)\approx m_{0}(\varrho^{2}-a^{2})^{3/2}$ and, to the fist order in $\rho^{2}-a^{2}$, the metric given by Eq. (11) reduces to
$ds^{2}_{e}=$ $\,-\left[1-2m_{0}\left(\varrho^{2}-a^{2}\right)\right]dt^{2}-4m_{0}a\left(\varrho^{2}-a^{2}\right)dtd\varphi$
$\;+d\varrho^{2}+\left[\varrho^{2}+2m_{0}a^{2}\left(\varrho^{2}-a^{2}\right)\right]d\varphi^{2},$ (12)
with $\varrho\in[a,\,a+\varepsilon)$ for $\varepsilon>0$.
In Ref. [7], the chosen exterior metric is
$ds^{2}_{e}=-\left[1-2m_{0}\left(\varrho^{2}-a^{2}\right)\right]dt^{2}+d\varrho^{2}+\varrho^{2}d\varphi^{2}.$ (13)
In this approximation as made in [7], the terms containing $m_{0}a$ and $m_{0}a^{2}$ that appear in (12) have been ignored and, as a consequence, the exterior metric results a static (nonrotating) metric. In fact, metric (13) may correspond to a de Sitter (dS) or an anti-de Sitter-like (AdS) region, depending on the sign of $m_{0}$, with the curvature scalars being constant. The authors in [7] have chosen $m_{0}>0$, what leads to a de Sitter-like region with positive curvature scalars.
III The matching at the ring
Even though the Darmois-Israel formalism (DIF) [27] for matching different metrics is well known, we review the main aspects of such a formalism here in order to set up notation and to present the important quantities adapted to the present problem.
Let us then refer the interior disk $\mathcal{D}$ with the metric given by Eq. (8) as the spacetime $(\mathcal{M}_{i},\,g_{i})$, and to the equatorial region outside the ring $\mathcal{A}$ with the metric given by Eq. (13) in a first study, and by (11) in a second study, as the spacetime $(\mathcal{M}_{e},\,g_{e})$. These two spacetimes are to be matched together at the ring $\mathcal{S}$. Let $\xi^{a}=(t,\varphi)$ be the intrinsic coordinates of the ring $\mathcal{S}$, whose metric is given by Eq. (9). Without lost of generality, we already identified the coordinates $t$ and $\varphi$ in the three regions $\mathcal{M}_{i}$, $\mathcal{M}_{e}$, and $\mathcal{S}$.
The DIF deals with the first and the second fundamental forms on $\mathcal{S}$, $h_{ab}$ and $K_{ab}$, respectively. Such quantities may be defined in terms of the geometric quantities in the spacetimes $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ by relations of the form $h_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}g_{\mu\nu}$ and $K_{ab}=e_{a}^{\ \mu}e_{b}^{\ \nu}\nabla_{\nu}n_{\mu}$, where $e_{a}^{\ \mu}=\partial x^{\mu}/\partial\xi^{a}$, $n_{\mu}$ is the spacelike unit normal vector to the ring $\mathcal{S}$, and $\nabla_{\nu}$ stands for the covariant derivative compatible with the Lorentzian metric. There is one of such relations for each one of the spacetimes $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$.
Moreover, in the DIF, it is necessary to choose the orientation of the unit normal vector to the boundary between the two spacetime regions, $n_{\mu}$, which we choose pointing from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$. The normal to a closed thin shell may have two distinct orientations relative to the center of the coordinates of each spacetime. In other words, the orientation of the normal with respect to the respective radial coordinate of each spacetime is not fixed a priory. Understanding this fact is essential for uncovering all possible thin shell solutions for the present problem.
The first boundary condition in the DIF is the continuity of the first fundamental form across the boundary $\mathcal{S}$, i.e., $[h_{ab}]=0$. Here we employ the standard notation $[Q]\equiv Q\big{|}^{e}_{\mathcal{S}}-Q\big{|}^{i}_{\mathcal{S}}$, with $Q\big{|}^{e}_{\mathcal{S}}$ denoting any given quantity $Q$ evaluated on $\mathcal{S}$ from the perspective of region $\mathcal{M}_{e}$, and similarly for $Q\big{|}^{i}_{\mathcal{S}}$. It is easy to see that, in the present case, the induced metric $h_{ab}$ at $\mathcal{S}$ as perceived from each spacetime is the same.
The second boundary condition establishes the relation between the possible discontinuity of the second fundamental form $K_{ab}$ across the boundary $\mathcal{S}$ and the energy-momentum (EMT) of the boundary layer. I.e., if $[K_{ab}]\neq 0$ then a thin shell at $\mathcal{S}$ is needed, and the respective energy-momentum tensor $S_{ab}$ is given by
$8\pi S_{ab}=-[K_{ab}]+h_{ab}[K],$ (14)
where $K=h^{ab}K_{ab}$.
IV Delving into the middle-Kerr: The nonrotating string
IV.1 The interior region
In the spacetime $(\mathcal{M}_{i},\,g_{i})$ interior to the ring, the metric is (9) and the spacelike unit normal vector to the boundary $\mathcal{S}$ as seen from $\mathcal{M}_{i}$ is given by
$n_{i\mu}=\epsilon_{i}\delta_{\mu}^{\rho},$ (15)
where $\epsilon_{i}$ is a normalization factor. The normalization condition $g^{\mu\nu}_{i}n_{i\mu}n_{i\nu}=1$ with metric (9) implies in $\epsilon_{i}=\pm 1$. The sign $\epsilon_{i}$ is the quantity that determines the relative orientation of $n_{i\mu}$ in relation to the coordinate $\rho$, as one can see from the scalar product between $n_{i\mu}$ and the gradient of the $\rho$, i.e., $\nabla_{\mu}\rho$. For $\epsilon_{i}=+1$, we have $g^{\mu\nu}_{i}n_{i\mu}(\nabla_{\nu}\rho)=+1>0$, which implies that the normal points in the direction of increasing $\rho$ as seen from $\mathcal{M}_{i}$. On the other hand, for $\epsilon_{i}=-1$, we have $g^{\mu\nu}_{i}n_{i\mu}(\nabla_{\nu}\rho)=-1<0$, which implies that the normal points in the direction of decreasing $\rho$ as seen from $\mathcal{M}_{i}$.
For $\rho=a$ the metric (8) defined in $\mathcal{M}_{i}$ reduces to the metric (9) defined in $\mathcal{S}$, and then the extrinsic curvature of $\mathcal{S}$ as seen from $\mathcal{M}_{i}$ is given by
$K^{t}_{it}=0,\ \ K^{t}_{i\varphi}=0=\ K^{\varphi}_{it},\ \ K^{\varphi}_{i\varphi}=\frac{\epsilon_{i}}{a}.$ (16)
IV.2 The exterior region
In the spacetime $(\mathcal{M}_{e},g_{e})$ exterior to the ring, the metric is (13) and the spacelike unit normal vector to the boundary $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is given by
$n_{\mu}^{+}=\epsilon_{e}\delta_{\mu}^{\varrho},$ (17)
where $\epsilon_{e}$ is the normalization factor. The condition $g^{\mu\nu}_{e}n_{e\mu}n_{e\nu}=1$ with metric (13) implies that $\epsilon_{e}=\pm 1$ and, therefore, determines the sign of $n_{e\mu}$. The sign $\epsilon_{e}$ is the quantity that determines the relative orientation between $n_{e\mu}$ and the coordinate $\varrho$. For $\epsilon_{e}=+1$, we have $g^{\mu\nu}_{e}n_{e\mu}(\nabla_{\nu}\varrho)=+1>0$, which implies that the normal points in the direction of increasing $\varrho$ as seen from $\mathcal{M}_{e}$. Whereas for $\epsilon_{i}=-1$, we have $g^{\mu\nu}_{e}n_{e\mu}(\nabla_{\nu}\varrho)=-1<0$, which implies that the normal points in the direction of decreasing $\varrho$ as seen from $\mathcal{M}_{e}$. Since, for $\varrho=a$, the metric (13) of $\mathcal{M}_{e}$ reduces to the metric (9) of $\mathcal{S}$, then the extrinsic curvature of $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is given by
$K^{t}_{et}=-2m_{0}a\epsilon_{e},\ \ K^{t}_{e\varphi}=0=\ K^{\varphi}_{et},\ \ K^{\varphi}_{e\varphi}=\frac{\epsilon_{e}}{a}.$ (18)
IV.3 Are all thin shell solutions possible?
IV.3.1 The matter-energy content at the ring
The Darmois-Israel formalism allows for both values of $\epsilon_{i}$ and $\epsilon_{e}$, which lead, in principle, to four possible thin shell solutions. However, the particular choices of $\epsilon_{i}$ and $\epsilon_{e}$ determine the possible domains of the coordinates $\rho$ and $\varrho$ which, at the end, have to be compared with the domains previously established for such coordinates.
[…]
$\sigma=-S^{t}_{\ t}=\frac{1}{8\pi a}(\epsilon_{i}-\epsilon_{e}),\ \ p=S^{\varphi}_{\ \varphi}=-\frac{m_{0}a}{4\pi}\epsilon_{e},$ (19)
are the energy-density and the pressure of the fluid, respectively, $h_{ab}$ is the metric of the ring given in Eq. (9), and $u_{a}=\delta_{a}^{\ t}$ is the fluid $2$-velocity at the ring. Hence, differently from the interpretation given in Ref. [7], we can see from the form of the fluid’s velocity that the string is not rotating.
Let us now look at the four possible different cases.
IV.3.2 The case $\epsilon_{i}=+1$ and $\epsilon_{e}=-1$
By starting with the choice made by Smailagic and Spallucci [7], i.e., $\epsilon_{i}=+1$ and $\epsilon_{e}=-1$, we get the linear energy density and the pressure of the string as $\sigma=1/4\pi a$ and $p=m_{0}a/4\pi$, respectively. The linear energy density is always positive, while the pressure can be positive or negative depending on the sign of $m_{0}$.
More importantly, let us notice that $\epsilon_{i}=+1$ implies that the normal vector points in the direction of increasing radial coordinate as seen from the interior region. Since the coordinate $\rho$ is in the interval $\rho\in[0,\,a)$, and recalling our convention that the normal vector $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the choice $\epsilon_{i}=+1$ implies that there is a center in the interior region, as expected. In turn, the choice $\epsilon_{e}=-1$ implies that the normal vector points in the direction of the decreasing radial coordinate as seen from the exterior region. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, this implies that in the exterior region one also has $\varrho\in[0,\,a]$. Hence, this possible solution has two centers, one in each region, which implies that we have a kind of Minkowski-dS or Minkowski-AdS closed universe that can be interpreted as a bubble universe, as similarly done in Ref. [29] for a Minkowski-Minkowski closed universe. However, the original domain of the coordinate $\varrho$ is assumed to be $\varrho\in[a,a+\varepsilon)$, for $\varepsilon>0$, with the already established notion that the exterior region corresponds to the equatorial plane just outside the ring. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
IV.3.3 The case $\epsilon_{i}=-1$ and $\epsilon_{e}=+1$
The other non trivial case is given by the choice $\epsilon_{i}=-1$ and $\epsilon_{e}=+1$. In this case, the linear energy density and the pressure of the fluid are given by $\sigma=-1/4\pi a$ and $p=-m_{0}a/4\pi$, respectively. The linear energy density is negative, while the pressure can be positive or negative depending on the sign of $m_{0}$.
More importantly, let us notice that $\epsilon_{i}=-1$ implies that the normal points in the direction of decreasing radial coordinate as seen from the interior region, while $\epsilon_{e}=+1$ implies that the normal points in the direction of increasing radial coordinate as seen from the exterior region. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, this implies that in both regions the radial coordinates start from $a$ and go to infinity. Therefore, the possible solution has no center at all and corresponds to a kind of Minkowski-dS or Minkowski-AdS open universe and can be interpreted as a traversable wormhole, as similarly done in Ref. [29] for a Minkowski-Minkowski open universe. However, once more let us notice that the original domain of the interior radial coordinate is $\rho\in[0,a]$, with the already established notion of the interior region corresponding to the region inside the ring, i.e., corresponding to the disk. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
IV.3.4 The case $\epsilon_{i}=-1$ and $\epsilon_{e}=-1$
For the choice $\epsilon_{i}=\epsilon_{e}=-1$, the linear energy density and the pressure of the fluid are given by $\sigma=0$ and $p=m_{0}a/4\pi$, respectively. The linear energy density is identically zero, while the pressure can be positive or negative depending on the sign of $m_{0}$.
In this case, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ are basically swapped. More importantly, let us notice that this choice $\epsilon_{i}$ and $\epsilon_{e}$ implies that the normal points in the direction of decreasing radial coordinate as seen from both regions. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the role of the radial coordinates $\rho$ and $\varrho$ are swapped. However, this contradicts the original domains of the radial coordinates and the notions that the interior region corresponds to the interior disk and the exterior region corresponds to the equatorial disk just outside the ring. Therefore, this choice does not lead to the appropriate solution for the thin shell at the boundary $\mathcal{S}$.
IV.3.5 The case $\epsilon_{i}=+1$ and $\epsilon_{e}=+1$
Finally, for the choice $\epsilon_{i}=\epsilon_{e}=+1$, the linear energy density and the pressure of fluid are given by $\sigma=0$ and $p=-m_{0}a/4\pi$. The linear energy density is identically zero, while the pressure can be positive or negative depending on the sign of $m_{0}$.
In this case, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ corresponds to the expected notion of interior and exterior regions. More importantly, let us notice that this choice for $\epsilon_{i}$ and $\epsilon_{e}$ implies that the normal points in the direction of increasing radial coordinate as seen from both regions. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the role of the radial coordinates $\rho$ and $\varrho$ stays the same with the domains given by $\rho\in[0,a]$ and $\varrho\in[a,a+\varepsilon)$ with $\varepsilon>0$, thus respecting the original domain of the radial coordinates. Therefore, this is the choice that corresponds to a consistent solution for the thin shell at $\mathcal{S}$, since it leads to a local geometry that is appropriate to match the global properties of the Kerr geometry with a regular core. Moreover, the usual assumption in order to construct regular black hole solutions is $m_{0}>0$, from which we notice that the strong energy condition is violated at the ring, what is in agreement with previous results [21].
V Delving into the middle-Kerr: The rotating string
V.1 The disk and the ring
The interior spacetime $(\mathcal{M}_{i},g_{i})$ is the disk $\mathcal{D}$ with the metric given by Eq. (8) as described in Sec. II.2. Hence, the spacelike unit normal vector to $\mathcal{S}$ as seen from $\mathcal{M}_{i}$ is the same as in the preceding case, given by (15), and, moreover, the extrinsic curvature of $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is also the same as in Eq. (16).
In turn, the ring $\mathcal{S}$ is described by the same metric given by Eq. (9).
V.2 The equatorial plane and the exterior metric
Let us remind that, as the ring is approached from the outside in the equatorial plane, the metric given by Eq. (11) reduces to the form (12). In contrast with the metric given by Eq. (13) used by the authors in Ref. [7], here we consider the full metric that comes from using $m(\varrho)\approx m_{0}(\varrho^{2}-a^{2})^{3/2}$ for $\varrho\to a$ in Eq. (11), without any further approximation. In other words, besides taking $\theta=\pi/2$ into the full metric, the only approximation we make is by choosing the asymptotic form of the mass function $m(\varrho)$.
Therefore, the exterior spacetime $(\mathcal{M}_{e},\,g_{e})$ now refers to the equatorial region outside the ring $\mathcal{A}$ with the metric given by Eq. (12). The spacelike unit normal vector to $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is given by (17), and, moreover, the extrinsic curvature of $\mathcal{S}$ is given by
$K^{t}_{et}=-2m_{0}a\epsilon_{e},\ \ K^{t}_{e\phi}=2m_{0}a^{2}\epsilon_{e},$
$K^{\varphi}_{et}=-2m_{0}\epsilon_{e},\ \ K^{\varphi}_{e\varphi}=\frac{\epsilon_{e}}{a}+2m_{0}a\epsilon_{e}.$ (20)
V.3 The rotating match at the ring
V.3.1 The matter-energy content at the ring
[…]
$8\pi S^{t}_{\ t}=2m_{0}a\epsilon_{e}+\frac{1}{a}(\epsilon_{e}-\epsilon_{i}),$ (21)
[…]
$8\pi S^{\varphi}_{\ \varphi}=-2m_{0}a\epsilon_{e}.$
Thus, we can see that, differently from the EMT given by Eq. (19), in this case the EMT of the shell is not diagonal. In fact, the presence of nondiagonal terms like $\mathcal{S}^{t}_{\varphi}$ is an indication that the fluid at the shell is rotating with respect to the asymptotic observer.
V.3.2 Orientations of the normal vector in which $\epsilon_{i}=-\epsilon_{e}$
If $\epsilon_{i}=-\epsilon_{e}$, we can diagonalize the EMT of the shell such that the eigenvalues are given by
$\lambda_{\pm}=\frac{\epsilon_{e}}{8\pi a}\left(1\pm\sqrt{1+4m_{0}a^{2}}\right),$ (22)
with the respectively normalized eigenvectors $v^{a}_{\pm}$ given by
$v^{t}_{\pm}=\sqrt{a^{2}\left(v^{\varphi}_{\pm}\right)^{2}\pm 1},$ (23)
$v^{\varphi}_{\pm}=\left[\pm\big{(}\omega^{-2}_{\pm}-a^{2}\big{)}\right]^{-1/2},$ (24)
where
$\omega_{\pm}$ $=\frac{S^{\varphi}_{\ t}}{\lambda_{\pm}-S^{\varphi}_{\ \varphi}}=\frac{2m_{0}a}{1+2m_{0}a^{2}\pm\sqrt{1+4m_{0}a^{2}}}.$ (25)
The timelike eigenvector that corresponds to the comoving observer with the shell is $v^{a}_{+}$, while the spacelike eigenvector is $v^{a}_{-}$. Hence, we can define the energy density of the shell as $\sigma=-\lambda_{+}=-S_{ab}v^{a}_{+}v^{b}_{+}$, while $p=\lambda_{-}=S_{ab}v^{a}_{-}v^{b}_{-}$ is the tangential pressure. Moreover, the eigenvector $v^{a}_{+}$ can be decomposed as
$v^{a}_{+}=\gamma\left(\delta^{a}_{\ t}+\omega_{+}\delta^{a}_{\ \varphi}\right),$ (26)
where $\gamma=\omega_{+}^{-1}/(\omega^{-2}_{+}-a^{2})^{1/2}$ and $\omega_{+}$ can be interpreted as the angular velocity of the rotating shell.
For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the energy density of the shell is positive, while the pressure can be either positive or negative depending on the sign of $m_{0}$. Similarly to the case discussed in Sec. IV.3.2, this scenario represents a closed universe with a rotating string at the ring joining the two regions.
For $\epsilon_{e}=1$, which also means $\epsilon_{i}=-1$, the energy density of the shell is negative, while the pressure can be either positive or negative depending on the sign of $m_{0}$. Moreover, similar to the case discussed in Sec. IV.3.3, this scenario corresponds to an open universe with a rotating string at the ring joining the two regions.
V.3.3 Orientations of the normal vector in which $\epsilon_{i}=\epsilon_{e}$
If $\epsilon_{i}=\epsilon_{e}$, it is not possible to diagonalize the EMT of the shell since the determinant of $S^{a}_{\ b}$ vanishes and, therefore, we are in the presence of a null dust fluid. This is indeed the case once that we can recast $S^{a}_{\ b}$ as
$S^{a}_{\ b}=Hk^{a}k_{b},$ (27)
where $k^{a}$ is a lightlike vector given by
$k^{a}=\delta^{a}_{\ t}+\frac{1}{a}\delta^{a}_{\ \varphi},$ (28)
with $\delta^{a}_{b}$ standing for the Kronecker delta. The quantity $H$ is given by
$H=-\frac{m_{0}a\epsilon_{e}}{4\pi}.$ (29)
In fact, in the present case it is a constant, and can be interpreted as the energy density of the shell. Hence, the thin-shell ring corresponds to a rotating string composed by a null dust fluid.
In the case with $\epsilon_{e}=1$, which also means $\epsilon_{i}=1$, the energy density depends on the sign of $m_{0}$, being positive for $m_{0}<0$. Notice that a negative $m_{0}$ means a negative mass function. Moreover, $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ correspond to the expected notion of interior and exterior regions, respectively, with both regions being joined by a rotating string at the ring. Hence, this is the appropriate choice to match the overall geometry of the Kerr-like spacetime with a regular rotating core. However, there is a drawback regarding the kind of matter of the string, the energy density of the string is negative for $m_{0}>0$. It is worth remembering that, when the extension through the region $r<0$ is not performed, the metric components of the Kerr spacetime are not smooth in the region $r=0$, what Israel [30] interpreted as caused by a layer of mass spread over the disk. The material distributed over the disk presents negative energy density and rotates with superluminal velocity. Here, the situation is not so extreme since the string is composed by a null dust with negative energy density. This implies that the null energy condition is violated at the ring.
In the case with $\epsilon_{e}=-1$, which also means $\epsilon_{i}=-1$, the energy density depends on the sign of $m_{0}$, being positive for $m_{0}>0$. Similarly to the case discussed in Sec. IV.3.4, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ are swapped and the resulting geometry does not correspond to the original problem.
VI Discussion
Following the proposal by Smailagic and Spallucci [7], in which the problematic central ring of the Kerr-like geometries is replaced with a string, we analyzed all possible thin shell configurations that can be built from the Darmois-Israel formalism regarding the possible orientations of the normal vector relative to the radial coordinate. The different orientations of the normal vector engender four different configurations and may be parameterized by two unit parameters $\epsilon_{e}=\pm 1$ and $\epsilon_{e}=\pm 1$.
Considering the original choice $\epsilon_{i}=-\epsilon_{e}=+1$ made by Smailagic and Spallucci [7], we conclude that the resulting solution is not appropriate for a thin shell at the ring of a rotating regular black hole, once the resulting spacetime topology contradicts the original domain of the exterior radial coordinate and the notion of the exterior region as the equatorial region just outside the ring. In fact, such a choice gives rise to a closed spacetime, that may be interpreted as a Minkowski-de Sitter closed universe, corresponding to a bubble universe.
More importantly, we formally demonstrated that the only consistent solution for the thin shell at the ring of a rotating regular black hole results from the choice $\epsilon_{i}=\epsilon_{e}=+1$. In this case, the original domains of the radial coordinates are respected and the roles of the interior and exterior regions are preserved. Hence, differently from Ref. [7], we find that the suitable string solution replacing the Kerr ring has a vanishing linear energy density and a pressure that can be either positive or negative depending on the sign of the constant $m_{0}$.
Finally, we noticed that in the scenario considered in Ref. [7], the string of matter replacing the ring is not rotating. In order to remedy this fact, we constructed a scenario where the string is indeed rotating, and considered all the four combinations of the normal vector orientations. The results for the corresponding geometries are the same as for the nonrotating string, while the kinds of material result quite different. In the cases where $\epsilon_{i}=-\epsilon_{e}$, the rotating string corresponds to a fluid with positive energy density for $\epsilon_{i}=-\epsilon_{e}=-1$, and with negative energy density for $\epsilon_{i}=-\epsilon_{e}=+1$. In the case where $\epsilon_{i}=\epsilon_{e}$, the rotating string corresponds to a null fluid with positive energy density for $m_{0}\epsilon_{e}<0$, and with negative energy density for $m_{0}\epsilon_{e}>0$.
Acknowledgements.
M. L. W. B. is funded by Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), Brazil, Grant No. 2022/09496-8. V. T. Z. thanks partial financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil, Grant No. 311726/2022-4, and from Fundação de Aperfeiçoamento do Pessoal de Nível Superior (CAPES), Brazil, Grant No. 88887.310351/2018-00.